OpenStudy (anonymous):

find the inverse of f(x)= (3^x)/(2-3^x)

OpenStudy (anonymous):

the x powers are the part that is confusing me, I know I have to solve for x and then switch the x and y... but I cant get to that poing

OpenStudy (lgbasallote):

first..change f(x) into y \[y = \frac{3^x}{2 - 3^x}\] cross multiply \[\implies y(2 - 3^x) = 3^x\] distribute \[\implies 2y - y3^x = 3^x\] put the x terms on one side \[\implies 2y = 3^x + y3^x\] factor out \[2 = 3^x(1 + y)\] divide both sides by 1 + y \[\frac{2}{1+y} = 3^x\] change to log form \[\log_3 \left (\frac{2}{1+y}\right) = x\] does that help?

OpenStudy (lgbasallote):

@cmgeorgia you there?

OpenStudy (anonymous):

yes, I was reviewing the problem. Thank you so much! One thing to clarify, shouldn't it be 2y/ 1+y)?

OpenStudy (lgbasallote):

ahh yes..good observation :D

OpenStudy (anonymous):

okay just making sure! Thanks again!

OpenStudy (lgbasallote):

welcome